# Calculation of RF-induced Temporal Jitter in Ultrafast Electron Diffraction

### v1

@@ I Introduction
Ultrafast electron diffraction (UED) has demonstrated great potential in visualizing the atmoic motions in different materials and molecules at the femto-second timescale [1]. UED employs a pump–probe scheme, where a pump laser pulse interacts with the sample and initiates the ultrafast process of interest, followed by a short electron bunch to probe the structural changes over different pump delays. Adjusting the delay between pump laser and probe electron allows time-resolved measurements of the structural evolution of the sample. The temporal resolution of a UED instrument is described by,
@@ II.1 Source of Timing Jitter
As we transition to a more quantitative description of jitters, the notation of all jitter quantities will be changed from $\delta$ to $d$ to represent the small deviation from nominal values as the later is conventionally used in differentiation. We also assume the difference of transport jitters in the pump and probe path is negligible ($\delta t_{\text{cathode}}-\delta t_{\text{pump}}\approx 0$). In practice, this assumption is valid when the laser transport path is short and the temperatures of all optical components are well regulated.
@@ III Timing Jitter Calculation
In this section we calculate timing jitter downstream of RF cavities from their phase and amplitude jitters. The longitudinal dynamics in a standing-wave RF cavity can be described as [2],
@@
We compare the timing jitters at exit ($z$=0.1 m) of the 1.5 cell S-band RF gun obtained by different methods. For benchmark, we generate a 10 fC electron beam distribution with 100 $\mu$m radius and 60 fs FWHM duration and track its time of arrival in the same RF gun field with gpt [14]. 100 simulations with different random realizations of the RF amplitude and phase were performed and the fractional amplitude and phase jitter values were randomly generated with a normal distribution with respective rms values of 500 ppm and 0.1 degree. As a benchmark, the magnitude of jitter values are larger than typically achieved values. As depicted in Figure 2, the analytical approximate solution overestimates the magnitude of timing jitters while the jitters calculated with semi-analytical model agree well with the gpt results (maxiumum deviation within 0.1 fs).
@@
Given the layout, cavity field maps, and operating parameters of a beamline, the numerical relation between RF and temporal jitter can be obtained from Eq. 15, with the Jacobian matrix in Eq. 18 and Eq. 20 calculated from an differentiable solver. The magnitude of temporal jitter can then be estimated based on performance of the klystron and laser-to-RF synchronization. In the following section, we’ll use the approach to study the impact of RF jitters in the SLAC MeV-UED instrument.
@@ IV Temporal Resolution of the SLAC MeV-UED Instrument
The SLAC MeV-UED instrument has been developing robust methods to generate ultrashort electron bunches with low arrival time jitters [6]. Currently, a LCLS-type 1.6 cell S-band photocathode rf gun serves as the electron source [16]. A 1.4 cell gun is being considered as the next generation electron source as it provides higher extraction field and bunch compression [17, 18].
@@
Here we compare the bunch length and timing jitter of the electron bunches produced from the 1.6 cell and 1.4 cell gun for different initial phases at $z=1.55$m (sample location). Both guns are assumed to operate at nominal gradients of 90 MV/m with $\alpha_{\mathrm{g}}=1.47$ The temporal jitters are calculated using our semi-analytical method, and the bunch length is modeled with particle tracking code gpt. For the jitter sources, we assume the fractional klystron voltage jitter to be 20 ppm (rms), and laser-to-rf synchronization jitter to be 25 fs (rms) [19]. The correlation coefficient between phase and klystron jitter $\mathcal{K}$ is -14.6 based on specification of the ScandiNova klystron system currently in use at the SLAC facility.
@@
Assuming negligible velocity mismatch between pump and probe, and including pump laser duration (25 fs rms), electron bunch length and timing jitter, the 1.6 cell gun achieves an overall temporal resolution of 46 fs (rms) when operating at 28 degree. The optimal phase for temporal resolution deviates from the phase of minimum jitter (around 40 degree) since the temporal resolution is dominated by the electron bunch length. For the 1.4 cell gun, the optimal temporal resolution is 43 fs when operating at 29 degree and the dominant contribution is the timing jitter.
@@ IV.2 Gun and Buncher
We consider a beamline with a 1.6 cell gun followed by a 5-cell S band buncher operating at 25 MV/m at $z=0.75$m from the cathode. We scan the initial phases for gun and buncher and perform particle-in-cell simulation to find the minimum bunch lengths for each initial phases and the corresponding waist locations. The timing jitters at waist locations are calculated using the klystron and synchronization parameters in Sec. IV.1, assuming the gun and buncher are powered by the same klystron—the fractional amplitude and phase jitters for the two cavities are identical. The calculation is then repeated for the scenario where the RF phase and amplitude jitters for gun and buncher are independent. The results of the minimum bunch lengths and the difference in timing jitters for the two cases are shown in Figure 5. For the same-klystron configuration, the rms timing jitters at waist do not show clear dependence on the initial phases of gun and buncher and are around 30 fs for the phase range considered. The dependence is more prominent for the different-klystron configuration and timing jitter is larger when operating at higher gun phases. The advantage of the same-klystron configuration only becomes apparent at higher gun phases.
@@
1.6 cell gun 37 9 46
1.4 cell gun 9 33 43
1.6 cell gun + buncher 5 30 40
1.4 cell gun + buncher 3 29 39
Table 1: RMS bunch length, arrival time jitter, and temporal resolution of different beamline configurations. The temporal resolution is calculated including the contribution of pump laser duration $\sigma_{\text{pump}}$=25 fs and assumes klystron voltage jitter of 20ppm (rms) and laser-to-rf synchronization jitter of 25fs (rms).
The optimal temporal resolution achieved by the 1.6 cell gun and the buncher is around 40 fs for the two configurations (gun and buncher powered by same or different klystrons). We also considered a beamline with the 1.4 cell gun and the buncher and found the level of arrival time jitter is similar. For brevity, the corresponding graphs are not included. The rms bunch duration, arrvial time jitter and temporal resolution of a 1.4 cell gun and buncher beamline are summarized in Table 1 along with other beamline configurations. Overall, the main contribution to temporal resolution in a gun-and-buncher beamline is the arrival time jitter and finite pump duration.
@@ IV.3 Jitter Correction
When the temporal resolution is dominated by the timing jitters, the approach outlined in Section III serves as a virtual timing tool to predict the RF-induced timing jitter. Since RF amplitude and phase jitters are collected every shot, the measured RF jitter values allows direct calculation of timing jitters with our approach. The predicted time jitter values can be used to sort the corresponding diffraction data and partially correct the arrival time jitters—the effect of laser-to-rf synchronization error remains and cannot be corrected.
We evaluate the correction approach by considering the timing jitters of the 1.4 cell gun beamline and the 1.6 cell gun and buncher beamline from previous sections. The 1.4 cell gun is assumed to operate at 28 deg, while the 1.6 cell gun and buncher operate at 30 deg phase and zero crossing respectively and are powered by different klystrons. We vary the level of klystron voltage jitter and calculate the rms timing jitter with and without correction. The rms value of laser-to-rf locking error is assumed to be 25fs. The results of the calculation are shown in Figure 6. The correction scheme reduces the jitters across different klystron jitter levels, with the residual jitter limited by laser-to-rf synchronization performance. To minimize the residual jitter, a THz streaking structure can be used to directly measure the time jitters relative to pump laser [22, 23, 24].
@@ VI Acknowledgements
This work was supported by the U.S. Department of Energy Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The SLAC MeV-UED program is supported by the U.S. Department of Energy Office of Science, Office of Basic Energy Science under FWP 10075, FWP 100713, and FWP 100940.

### v2

@@ I Introduction
Ultrafast electron diffraction (UED) has demonstrated great potential in visualizing the atomic motions in different materials and molecules at the femtosecond timescale [1]. UED employs a pump–probe scheme, where a pump laser pulse interacts with the sample and initiates the ultrafast process of interest, followed by a short electron bunch to probe the structural changes over different pump delays. Adjusting the delay between pump laser and probe electron allows time-resolved measurements of the structural evolution of the sample. The temporal resolution of a UED instrument is described by,
@@ II.1 Source of Timing Jitter
As we transition to a more quantitative description of jitters, the notation of all jitter quantities will be changed from $\delta$ to $d$ to represent the small deviation from nominal values as the latter is conventionally used in differentiation. We also assume the difference of transport jitters in the pump and probe path is negligible ($\delta t_{\text{cathode}}-\delta t_{\text{pump}}\approx 0$). In practice, this assumption is valid when the laser transport path is short and the temperatures of all optical components are well regulated.
@@ II.2 Correlated Phase and Amplitude Jitter
In the presence of an active feedback system, the amplitude and phase jitters of the RF gun are reduced from the values in Eq. 6 and Eq. 10. However, the reduction is usually not perfect and the residual jitters still contribute to timing jitters.
@@ III Timing Jitter Calculation
In this section we calculate timing jitters downstream of RF cavities from their phase and amplitude jitters. The longitudinal dynamics in a standing-wave RF cavity can be described as [2],
@@
We compare the timing jitters at exit ($z$=0.1 m) of the 1.5 cell S-band RF gun obtained by different methods. For benchmark, we generate a 10 fC electron beam distribution with 100 $\mu$m radius and 60 fs FWHM duration and track its time of arrival in the same RF gun field with gpt [14]. 100 simulations with different random realizations of the RF amplitude and phase were performed and the fractional amplitude and phase jitter values were randomly generated with a normal distribution with respective rms values of 500 ppm and 0.1 degree. As a benchmark, the magnitude of jitter values are larger than typically achieved values. As depicted in Figure 2, the analytical approximate solution overestimates the magnitude of timing jitters while the jitters calculated with semi-analytical model agree well with the gpt results (maximum deviation within 0.1 fs).
@@
Given the layout, cavity field maps, and operating parameters of a beamline, the numerical relation between RF and temporal jitter can be obtained from Eq. 15, with the Jacobian matrix in Eq. 18 and Eq. 20 calculated from a differentiable solver. The magnitude of temporal jitter can then be estimated based on performance of the klystron and laser-to-RF synchronization. In the following section, we’ll use the approach to study the impact of RF jitters in the SLAC MeV-UED instrument.
@@ IV Temporal Resolution of the SLAC MeV-UED Instrument
The SLAC MeV-UED instrument has been developing robust methods to generate ultrashort electron bunches with low arrival time jitters [6]. Currently, an LCLS-type 1.6 cell S-band photocathode rf gun serves as the electron source [16]. A 1.4 cell gun is being considered as the next generation electron source as it provides higher extraction field and bunch compression [17, 18].
@@
Here we compare the bunch length and timing jitter of the electron bunches produced from the 1.6 cell and 1.4 cell gun for different initial phases at $z=1.55$m (sample location). Both guns are assumed to operate at nominal gradients of 90 MV/m with $\alpha_{\mathrm{g}}=1.47$ The timing jitters are calculated using our semi-analytical method, and the bunch length is modeled with particle tracking code gpt. For the jitter sources, we assume the fractional klystron voltage jitter to be 20 ppm (rms), and laser-to-rf synchronization jitter to be 10 fs (rms) [19]. The correlation coefficient between phase and klystron jitter $\mathcal{K}$ is -14.6 based on specification of the ScandiNova klystron system currently in use at the SLAC facility. The effects of RF feedback systems are not included in the jitter calculation and therefore the jitter values should be slightly overestimated.
@@
Assuming negligible velocity mismatch between pump and probe, and including pump laser duration (25 fs rms), electron bunch length and timing jitter, the 1.6 cell gun achieves an overall temporal resolution of 46 fs (rms) when operating at 25 degree. The optimal phase for temporal resolution deviates from the phase of minimum jitter (around 40 degree) since the temporal resolution is dominated by the electron bunch length. For the 1.4 cell gun, the optimal temporal resolution is 33 fs when operating at 27 degree and the dominant contribution is the timing jitter and pump laser duration.
@@ IV.2 Gun and Buncher
We consider a beamline with a 1.6 cell gun followed by a 5-cell S band buncher operating at 25 MV/m at $z=0.75$m from the cathode. We scan the initial phases for gun and buncher and perform particle-in-cell simulation to find the minimum bunch lengths for each initial phases and the corresponding waist locations. The timing jitters at waist locations are calculated using the klystron and synchronization parameters in Sec. IV.1, assuming the gun and buncher are powered by the same klystron—the fractional amplitude and phase jitters for the two cavities are identical. The calculation is then repeated for the scenario where the RF phase and amplitude jitters for gun and buncher are independent. The results of the minimum bunch lengths and the difference in timing jitters for the two cases are shown in Figure 5. For the same-klystron configuration, the rms timing jitters at waist do not show clear dependence on the initial phases of gun and buncher and are around 18 fs for the phase range considered. The dependence is more prominent for the different-klystron configuration and timing jitter is larger when operating at higher gun phases. The advantage of the same-klystron configuration only becomes apparent at higher gun phases.
@@
1.6 cell gun 37 6 46
1.4 cell gun 7 20 33
1.6 cell gun + buncher 6 18 31
1.4 cell gun + buncher 4 18 31
Table 1: RMS bunch length, arrival time jitter, and temporal resolution of different beamline configurations. The temporal resolution is calculated including the contribution of pump laser duration $\sigma_{\text{pump}}$=25 fs and assumes klystron voltage jitter of 20ppm (rms) and laser-to-rf synchronization jitter of 10fs (rms).
The optimal temporal resolution achieved by the 1.6 cell gun and the buncher is around 31 fs for the two configurations (gun and buncher powered by same or different klystrons). We also considered a beamline with the 1.4 cell gun and the buncher and found the level of arrival time jitter is similar. For brevity, the corresponding graphs are not included. The rms bunch duration, arrvial time jitter and temporal resolution of a 1.4 cell gun and buncher beamline are summarized in Table 1 along with other beamline configurations. Overall, the main contribution to temporal resolution in a gun-and-buncher beamline is the arrival time jitter and finite pump duration.
@@ IV.3 Jitter Correction
When the temporal resolution is dominated by the timing jitters, the approach outlined in Section III serves as a virtual timing tool to predict the RF-induced timing jitter. Since RF amplitude and phase jitters are collected every shot, the measured RF jitter values allow direct calculation of timing jitters with our approach. The predicted time jitter values can be used to sort the corresponding diffraction data and partially correct the arrival time jitters—the effect of laser-to-rf synchronization error remains and cannot be corrected.
We evaluate the correction approach by considering the timing jitters of the 1.4 cell gun beamline and the 1.6 cell gun and buncher beamline from previous sections. The 1.4 cell gun is assumed to operate at 26 deg, while the 1.6 cell gun and buncher operate at 30 deg phase and zero crossing respectively and are powered by different klystrons. We vary the level of klystron voltage jitter and calculate the rms timing jitter with and without correction. The rms value of laser-to-rf locking error is assumed to be 10 fs. The results of the calculation are shown in Figure 6. The correction scheme reduces the jitters across different klystron jitter levels, with the residual jitter limited by laser-to-rf synchronization performance. To minimize the residual jitter, a THz streaking structure can be used to directly measure the time jitters relative to pump laser [22, 23, 24].
@@ VI Acknowledgements
We thank Lili Ma (SLAC) for useful discussion on laser-to-rf synchronization. This work was supported by the U.S. Department of Energy Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The SLAC MeV-UED program is supported by the U.S. Department of Energy Office of Science, Office of Basic Energy Science under FWP 10075, FWP 100713, and FWP 100940.